Ph calculation of 5 ml of 0.1 M BOhand 25ml of 0.1 M Bcl

Grade:11

1 Answers

BIKRAM
6 years ago
BOH = B+ + OH- : BCl = B+ + Cl- :total volume of solution = (5 +25 )= 30 ml;
Only OH- ion contributes towards PH calculation.[OH-]= (0.1× 5)/30 =1/60 (M)
So, [H+]×[OH-]=10-14; [H+] =60×10-14 ;PH= -log[H+] =12.22